# CLQCD Collaboration

### v2

@@ Abstract
We present the determination of the charm quark mass, the masses and decay constants of charmed mesons using thirteen 2+1 flavor full-QCD gauge ensembles at five different lattice spacings $a\in[0.05,0.11]$ fm, 8 pion masses $m_{\pi}\in(130,360)$ MeV, and several values of the strange quark mass, which facilitate us to do the chiral and continuum extrapolation. These ensembles are generated through the stout smeared clover fermion action and Symanzik gauge actions with the tadpole improvement. Using QED-subtracted $D_{s}$ meson mass and non-perturbative renormalization, we predict the charm quark mass in the continuum with physical light and strange quark masses to be $m_{c}(m_{c})=1.289(17)$ GeV in $\overline{\textrm{MS}}$ scheme, with the systematic uncertainties from lattice spacing determination, renormalization constant, and fit ansatz included. Predictions of the open and close charm mesons using this charm quark mass agree with the experimental value at 0.3% level uncertainty. We obtained $D_{(s)}$ decay constants and also by far the most precise $D_{(s)}^{*}$ decay constants $f_{D^{*}}=0.2321(43)$ GeV and $f_{D^{*}_{s}}=0.2743(34)$ GeV, with the charm quark improved vector current normalization.
@@ I Introduction
Lattice QCD has proven to be highly effective in providing first-principle predictions for the non-perturbative characteristics of light quarks with $m_{q}\ll\Lambda_{\rm QCD}$. On the other hand, the lattice QCD prediction of the hadron spectrum and structure involving heavy quarks with $m_{Q}\gg\Lambda_{\rm QCD}$ is also highly demanded. This is necessary for constraining the CKM matrix elements through the weak decay constant of hadrons, as well as for determining Higgs coupling to heavy quarks. Furthermore, lattice QCD calculations are crucial for revealing the nature of extraordinary hadron states such as tetraquarks Ablikim et al. (2013); Liu et al. (2013) and pentaquarks Wu et al. (2010); Aaij et al. (2015).
However, the lattice QCD calculations for heavy quarks with $m_{Q}\gg\Lambda_{\rm QCD}$ are significantly more challenging. The primary difficulty arises from the fact that the heavy quark mass can introduce additional discretization errors of the order $m_{Q}^{2}a^{2}$, and even the $m_{Q}^{4}a^{4}$ corrections can become significant when $a\sim 0.1$ fm and $m_{Q}\sim 1$ GeV. Thus comparing to the light flavor case, the calculations at more lattice spacings are very beneficial to control the systematic uncertainty from the discretization error.
In this work, we use the 2+1 flavor full-QCD ensembles generated by the CLQCD collaborations at 5 lattice spacings to determine the charm quark mass and charmed meson properties. As discussed in the recent lattice average of flavor physics Aoki et al. (2022), the systematic uncertainty stemming from the absence of the charm quark as a dynamical degree of freedom in the action is expected to be at a level of a few percent. This is due to the decoupling of the heavy quark with $m_{Q}\gg\Lambda_{\rm QCD}$, and will be revisited once dynamical charm ensembles become available. Additionally, the QED effects of the charm quark mass are considered based on calculations from the literature Giusti et al. (2017), although they are not included in our predictions of hadron masses and decay constants. We expect that this systematic uncertainty will be under control in the near future, using the infinite volume reconstruction techniques Feng and Jin (2019).
Comparing to the previous charm quark mass determination using the clover fermion action Heitger et al. (2021), our calculation applies the stout smearing on the gauge field used by the fermion action, thereby offering a validation of the systematic uncertainty associated with the fermion action definition. We also investigate the sensitivity of meson decay constants to the discretization of the normalization constant $Z_{V}$, noting that this sensitivity is eliminated by the fact that the renormalization of the PCAC quark mass only requires the ratio $Z_{A}/Z_{P}$ of the renormalization constants.
@@
The organization of this work is outlined as follows. Section II presents the numerical setup, which includes the interpolation of strange and charm quark masses, the definition of different normalization constants, and their lattice spacing dependence. Section III shows the continuum extrapolation of the renormalized charm mass and hadron properties, and also comparison with the experiment and previous calculations. Finally, Section IV offers a concise summary and discussion.
II Numerical setup
@@
Table 1: Gauge coupling $\hat{\beta}=10/(g^{2}u_{0}^{4})$, lattice spacing $a$, tadpole improvement factors $u_{0}$ and $v_{0}$, dimensionless bare quark mass parameters $\tilde{m}^{\rm b}_{l,s}$, Lattice size $\tilde{L}^{3}\times\tilde{T}$, corresponding pion mass $m_{\pi}$ and $\eta_{s}$ mass $m_{\eta_{s}}$, interpolated bare strange and charm quark mass parameters $\tilde{m}^{\rm I}_{s,c}$, and the statistics information.
Symbol $\hat{\beta}$ $a$ (fm) $u_{0}$ $v_{0}$ $\tilde{m}^{b}_{l}$ $\tilde{m}^{b}_{s}$ $\tilde{L}^{3}\times\tilde{T}$ $m_{\pi}$ (MeV) $m_{\eta_{s}}$ (MeV) $\tilde{m}^{\rm I}_{s}$ $\tilde{m}^{\rm I}_{c}$ $n_{\rm cfg}\times n_{\rm src}$
C24P34 6.200 0.10530(18) 0.855453 0.951479 -0.2770 -0.2310 $24^{3}\times 64$ 340.5(1.7) 748.99(73) -0.2396(2) 0.4080(26) $200\times 32$
C24P29 0.855453 0.951479 -0.2770 -0.2400 $24^{3}\times 72$ 292.7(1.2) 658.29(65) -0.2357(2) 0.4168(26) $760\times\ \ \!3$
C32P29 0.855453 0.951479 -0.2770 -0.2400 $32^{3}\times 64$ 292.4(1.1) 659.22(41) -0.2358(2) 0.4158(26) $489\times\ \ \!3$
C32P23 0.855520 0.951545 -0.2790 -0.2400 $32^{3}\times 64$ 228.0(1.2) 644.36(45) -0.2338(2) 0.4198(26) $400\times\ \ \!3$
C48P23 0.855520 0.951545 -0.2790 -0.2400 $48^{3}\times 96$ 225.6(0.9) 644.58(62) -0.2338(2) 0.4214(26) $\ \ \!62\times\ \ \!3$
C48P14 0.855548 0.951570 -0.2825 -0.2310 $48^{3}\times 96$ 135.5(1.6) 707.06(44) -0.2335(2) 0.4212(26) $188\times\ \ \!3$
E28P35 6.308 0.08877(30) 0.859646 0.954385 -0.2490 -0.2170 $28^{3}\times 64$ 352.1(1.2) 720.31(94) -0.2204(3) 0.2707(37) $147\times\ \ \!4$
F32P30 6.410 0.07750(18) 0.863437 0.956942 -0.2295 -0.2050 $32^{3}\times 96$ 303.2(1.3) 677.6(1.0) -0.2039(2) 0.1968(21) $250\times\ \ \!3$
F48P30 0.863473 0.956984 -0.2295 -0.2050 $48^{3}\times 96$ 303.4(0.9) 676.32(62) -0.2038(2) 0.1949(21) $\ \ \!99\times\ \ \!3$
F32P21 0.863488 0.957017 -0.2320 -0.2050 $32^{3}\times 64$ 210.9(2.2) 660.27(94) -0.2024(2) 0.1989(21) $194\times\ \ \!3$
F48P21 0.863499 0.957006 -0.2320 -0.2050 $48^{3}\times 96$ 207.2(1.1) 663.39(65) -0.2026(2) 0.1991(21) $\ \ \!82\times 12$
G36P29 6.498 0.06826(27) 0.866476 0.958910 -0.2150 -0.1926 $36^{3}\times 108$ 295.1(1.2) 693.2(1.0) -0.1929(2) 0.1378(28) $\ \ \!68\times\ \ \!4$
H48P32 6.720 0.05187(26) 0.873378 0.963137 -0.1850 -0.1700 $48^{3}\times 144$ 317.2(0.9) 695.9(1.3) -0.1703(2) 0.0533(24) $157\times 12$
The results in this work, are based on the 2+1 flavor full QCD ensembles from the CLQCD collaboration using the tadpole improved tree level Symanzik (TITLS) gauge action and the tadpole improved tree level Clover (TITLC) fermion action. In addition to the ensembles employed for determining the light quark masses and low energy constants Hu et al. (2024), we have included two additional ensembles, E28P35 and G36P29, with $m_{\pi}\sim 300$ MeV and gauge couplings $\hat{\beta}=6.308$ and 6.498 to manage discretization errors. By using these two values of $\hat{\beta}$, self consistent $u_{0}$ and $v_{0}$, and also the lattice spacing determined through the gradient flow Lüscher (2010) with $w_{0}$ Borsanyi et al. (2012) using the Symanzik action, the spatial volume of the E28P35 and G36P29 ensembles are similar to those of the other ensembles with $m_{\pi}\sim$ 300 MeV. The parameters employed for the simulation are outlined in Table 1, and we use $\tilde{O}$ for the dimensionless value of any quantity $O$.
The variance of binned means of the plaquette $\square=\left\langle\operatorname{Tr}U_{p}\right\rangle/3$, topological charge $Q$, pion mass $m_{\pi}$, and $\eta_{c}$ mass $m_{\eta_{c}}$ as shown in the appendix, suggest that the autocorrelation between configurations is acceptable. This is particularly evident in the case of $Q$, which does not exhibit any indications of topological charge freezing at the smallest lattice spacing $a=0.0519(3)$ fm.
@@
Figure 1: The $m_{\eta_{s}}^{2}$ dependence of the $D_{s}^{*}$ mass at $a=0.1053$ fm, with unitary strange quark mass parameter (red dots) and interpolated one (blue crosses).
@@
Given that the unitary strange quark mass on some ensembles deviate by 10-15% from the physical strange quark mass, we have adjusted the interpolated valence strange quark mass $\tilde{m}^{\rm I}_{s}$ to that correspond to the “physical” $\eta_{s}$ mass of 689.89(49) MeV Borsanyi et al. (2021) on each ensemble to mitigate this discrepancy. This $\eta_{s}$ mass aligns with our previous determination $m_{\eta_{s}}=687.4(2.2)$ MeV on CLQCD ensembles but with significantly reduced uncertainties. Fig. 1 illustrates the strange quark mass dependence (through $m_{\eta_{s}}^{2}$) of the $D_{s}^{*}$ meson mass at $a=0.1053$ fm and interpolated charm quark mass, using either the unitary strange quark mass parameter $\tilde{m}_{s}$ (red circles) or the interpolated parameter $\tilde{m}^{\rm I}_{s}$ (blue crosses). The strange quark mass dependence of the latter case is only 10% of the former case and then allows a more reliable continuum extrapolation since the contribution of the cross term likes $m_{\eta_{s}}^{2}a^{2}$ will be highly suppressed. This residual mismatch effect of the sea strange quark mass will be addressed through a joint fit using all the ensembles.
@@
Determining the valence charm quark mass can present additional complexities. The physical $J/\psi$ or $\eta_{c}$ mass should account for both the disconnected charm quark contribution and QED effect, making it unsuitable for accurately determining the charm quark mass using pure QCD calculations with only the connected charm quark contribution. Conversely, the open-charm pseudoscalar meson masses is free of the disconnected valence charm quark contribution, and its QED effect has been computed in previous literature Giusti et al. (2017) and updated in Ref. Di Carlo et al. (2019). Compared to the strange quark, the valence light quark with physical mass is noisy, costly, and also susceptible to significant finite-volume effects on most of our ensembles. Thus the optimal choice to determine the interpolated valence charm quark mass parameter $\tilde{m}^{\rm I}_{c}$ on each ensemble, is requiring the $D_{s}$ meson mass to be the QED effect subtracted physical value Di Carlo et al. (2019) under the GRS renormalization scheme $m^{\overline{\textrm{MS}}}_{q,\rm QCD+QED}(2\mathrm{GeV})=m^{\overline{\textrm{MS}}}_{q,\rm QCD}(2\mathrm{GeV})$ Gasser et al. (2003),
@@
Figure 2: The ratio of the PCAC quark mass determined from the quarkoinum, and also $D_{s}$. Two definitions are consistent up to ${\cal O}(a^{4})$ correction.
The values of the interpolated strange and charm quark masses $\tilde{m}^{\rm I}_{s,c}$ are also presented in Table 1, based on the polynomial interpolation of the calculations of $m_{\eta_{s},D_{s}}$ using three strange (charm) mass parameters. The sensitivity of $\tilde{m}^{\rm I}_{s,c}$ to the light and strange sea quark mass and lattice volume is on the order of a few MeV and should be eliminated upon extrapolation to physical quark masses and the infinite volume limit.
@@
where $A_{\mu}=\bar{\psi}\gamma_{5}\gamma_{\mu}\psi$, $P=\bar{\psi}\gamma_{5}\psi$, and $m_{\rm PS}$ is the corresponding pseudoscalar meson mass.
In Fig. 2, we show the ratio of two determinations of $m_{s}+m_{c}$ from quarkonium ($\eta_{s}$ and $\eta_{c}$) and also $D_{s}$, at five lattice spacings, with the interpolated strange and charm quark masses. Two determinations deviate from each other by approximately 1.4% at the coarsest lattice spacing but are consistent within the statistical uncertainty of 0.1% after a polynomial extrapolation up to $\mathcal{O}(a^{4})$ to the continuum limit.
The renormalized quark mass can be subsequently defined as $m^{R}_{q}=Z_{A}/Z_{P}m^{\rm PC}_{q}$. The ratio $Z_{A}/Z_{P}$ can be extracted from the corresponding amputated vertex functions $\Lambda_{A,P}$ in the off-shell quark state with momentum $p$ under the Landau gauge Hu et al. (2024),
subtracting the $1/m_{q}$ mass pole of the Goldstone meson Liu et al. (2014) and extrapolating to the chiral limit. $\Lambda_{\cal O}(p)=S^{-1}(p)G_{\cal O}(p)S^{-1}(p)$ can be obtained from the quark propagators $S(p)=\sum_{x}e^{-ip\cdot x}\langle\psi(x)\bar{\psi}(0)\rangle$ and also $G_{\mathcal{O}}(p)=\sum_{x,y}e^{-ip\cdot(x-y)}\langle\psi(x)\mathcal{O}(0)\bar{\psi}(y)\rangle$.
Following the previous light quark study on the same ensembles, we use the coulomb wall source propagator to suppress the statistical uncertainty, and extract the meson mass and local operator matrix elements through the joint fit of the wall-to-point and wall-to-wall two point functions Hu et al. (2024). The number of the configuration $n_{\rm cfg}$ and sources used per configuration $n_{\rm src}$ are also collect in Table 1. Then the meson decay constants can be extracted from the matrix elements with proper normalization,
@@
For the clover fermion, the ratio between $Z_{A}$ and $Z_{V}$,
$\frac{Z_{A}}{Z_{V}}$ $=\frac{\frac{1}{12}\text{Tr}[\Lambda_{V}(p)\gamma_{\mu}]}{\frac{1}{48}\text{Tr}[\Lambda^{\mu}_{A}(p)\gamma_{\mu}\gamma_{5}]},$ (6)
can deviate from 1 due to additive chiral symmetry breaking, and it is essential to include this effect to obtain correct $f_{\pi,K}$ after the continuum extrapolation Hu et al. (2024). More detailed discussion on the ratio
$\frac{Z_{T}}{Z_{V}}$ $=\frac{\frac{1}{12}\text{Tr}[\Lambda_{V}(p)\gamma_{\mu}]}{\frac{1}{72}\text{Tr}[\Lambda^{\mu\nu}_{T}(p)\gamma_{\nu}\gamma_{\mu}]},$ (7)
can be found in our previous work Hu et al. (2024).
Figure 3: The dimensionless pseudo-scalar mass $m_{\rm PS}^{2}a^{2}$ dependence of $Z_{V}$ using the pseudo-scalar meson mass, at different lattice spacing.
Therefore, all renormalization constants can be obtained by taking their ratio over $Z_{V}$, once $Z_{V}$ is determined. But unlike its continuum counterpart, $Z_{V}$ under lattice regularization is subject to both $\alpha_{s}$ and $a^{2}$ corrections and can be determined from the vector current conservation condition Martinelli et al. (1995); Zhang et al. (2021),

### v3

@@ Abstract
We present the determination of the charm quark mass, the masses, and decay constants of charmed mesons using thirteen 2+1 flavor gauge ensembles at five different lattice spacings $a\in[0.05,0.11]$ fm, 8 pion masses $m_{\pi}\in(130,360)$ MeV, and several values of the strange quark mass, which facilitate us to do the chiral and continuum extrapolation. These ensembles are generated through the stout smeared clover fermion action and Symanzik gauge actions with the tadpole improvement. By absorbing the discretization errors into the masses and field normalization of the charm quark, we manage to suppress the discretization error of the charmed meson mass and all the S-wave open charmed meson decay constants to a few percent or even less at lattice spacing $a\sim 0.1$ fm. Moreover, discretization errors for other quantities are also significantly reduced. The continuum extrapolated charm quark mass, $m_{c}(m_{c})=1.2933(72)(95)$ GeV in $\overline{\textrm{MS}}$ scheme, is determined using QED-subtracted $D_{s}$ meson mass and non-perturbative renormalization. Predictions of the open and close charm mesons using this charm quark mass agree with the experimental values at 0.1-0.5% level uncertainty. We obtained $D_{(s)}$ decay constants and also by far the most precise $D_{(s)}^{*}$ decay constants $f_{D^{*}}=0.2292(26)(17)$ GeV and $f_{D^{*}_{s}}=0.2691(30)(03)$ GeV.
@@ I Introduction
Lattice QCD has proven to be highly effective in providing first-principle predictions for the non-perturbative characteristics of light quarks with $m_{q}\ll\Lambda_{\rm QCD}$. On the other hand, the lattice QCD prediction of the hadron spectrum and structure involving heavy quarks with $m_{Q}\gg\Lambda_{\rm QCD}$ is also highly demanded. This is necessary for constraining the CKM matrix elements through the weak decay constant of hadrons, as well as for determining Higgs coupling to heavy quarks. Furthermore, lattice QCD calculations are crucial for revealing the nature of extraordinary hadron states such as tetraquarks Ablikim et al. (2013); Liu et al. (2013) and pentaquarks Wu et al. (2010); Aaij et al. (2015, 2019).
However, the lattice QCD calculations for heavy quarks with $m_{Q}\gg\Lambda_{\rm QCD}$ are significantly more challenging. As an effective theory of QCD with an intrinsic cutoff of the order of the inverse lattice spacing $1/a$, lattice QCD is subject to power corrections of the form $\left[\frac{\Lambda_{\rm QCD}}{1/a}\right]^{2}=\Lambda_{\rm QCD}^{2}a^{2}$. These corrections, which can also be viewed as discretization errors, can be eliminated by the continuum extrapolation $a\rightarrow 0$ (or equivalently $1/a\rightarrow\infty$). In the context of heavy quark physics, additional power corrections arise at ${\cal O}(m_{Q}^{2}a^{2})$ for most of the discretized fermion actions. Even the $m_{Q}^{4}a^{4}$ corrections can become significant when $a\sim 0.1$ fm and $m_{Q}\sim 1$ GeV, based on a naive power counting. Since performing lattice calculations with finer spacing and physical light quark masses demands substantially more computational resources than calculations using either heavier light quark masses or coarser lattice spacing, the precision calculation of the heavy flavor physics can be much more costly compared to the light flavor physics.
Similar to other effective theories, the convergence of the ${\cal O}(m_{Q}^{2}a^{2})$ corrections can be improved by redefining the expansion parameter $m_{Q}$ at a given cutoff of $1/a$, analogous to replacing $F_{\pi}(0)$ with $F_{\pi}(m_{\pi})$ in chiral perturbation theory. Moreover, most of the $m_{Q}a$ dependence to the charmed hadron matrix elements can be absorbed through the redefinition of the charm quark field at finite $1/a$, as we will demonstrate in this work. These improvements enable us to explore heavy quark physics, including its dependence on light quark mass, at relatively small values of $1/a$ with controllable power correction. Additionally, one can estimate residual power corrections by using ensembles at different lattice spacings but heavier light quark masses.
In this work, we conduct a systematic study of how charm physics observables, including the masses of open and hidden charm mesons as well as their decay constants, depend on the lattice spacing $a$ after implementing the improvements mentioned above. At $a\sim 0.1$ fm, the charmed meson mass and all the S-wave open charmed meson decay constants deviate from their continuum extrapolated values by only a few percent or even less, and discretization errors for other quantities are also significantly reduced. Although the redefined charm quark mass is subject to substantial ${\cal O}(a^{4})$ errors, its continuum extrapolated value remains consistent with current lattice averages Aoki et al. (2022), exhibiting an 1% uncertainty.
@@
The organization of this work is outlined as follows. Section II presents the theoretical framework, which includes the numerical setup, redefinition of the quark mass, and also renormalization of the matrix elements with the redefined charm quark field. Section III shows the continuum extrapolation of the renormalized charm mass and hadron properties, and also a comparison with the experiment and previous calculations. Finally, Section IV offers a concise summary and discussion.
II Theoretical framework
@@
Table 1: Gauge coupling $\hat{\beta}=10/(g^{2}u_{0}^{4})$, lattice spacing $a$ with the second error from that of the scale parameter $w_{0}$, dimensionless bare quark mass parameters $\tilde{m}^{\rm b}_{l,s}$, Lattice size $\tilde{L}^{3}\times\tilde{T}$, corresponding pion mass $m_{\pi}$ and $\eta_{s}$ mass $m_{\eta_{s}}$, the value of $m_{\pi}L$, valence bare strange and charm quark mass parameters $\tilde{m}^{\rm v}_{s,c}$, and the statistics information.
Symbol $\hat{\beta}$ $a$ (fm) $\tilde{m}^{\rm b}_{l}$ $\tilde{m}^{\rm b}_{s}$ $\tilde{L}^{3}\times\tilde{T}$ $m_{\pi}$ (MeV) $m_{\eta_{s}}$ (MeV) $m_{\pi}L$ $\tilde{m}^{\rm v}_{s}$ $\tilde{m}^{\rm v}_{c}$ $n_{\rm cfg}\times n_{\rm src}$
C24P34 6.200 0.10524(05)(62) --0.2770 --0.2310 $24^{3}\times 64$ 340.2(1.7) 748.61(75) 4.360(22) --0.2396(1) 0.4072(07) $200\times 32$
C24P29 --0.2770 --0.2400 $24^{3}\times 72$ 292.3(1.0) 657.83(64) 3.746(13) --0.2356(1) 0.4159(07) $760\times\ \ \!3$
C32P29 --0.2770 --0.2400 $32^{3}\times 64$ 293.1(0.8) 658.80(43) 5.008(14) --0.2358(1) 0.4150(06) $489\times\ \ \!3$
C32P23 --0.2790 --0.2400 $32^{3}\times 64$ 227.9(1.2) 643.93(45) 3.894(20) --0.2337(1) 0.4190(07) $400\times\ \ \!3$
C48P23 --0.2790 --0.2400 $48^{3}\times 96$ 224.1(1.2) 644.08(62) 5.743(30) --0.2338(1) 0.4196(08) $\ \ \!62\times\ \ \!3$
C48P14 --0.2825 --0.2310 $48^{3}\times 96$ 136.4(1.7) 706.55(39) 3.495(44) --0.2335(1) 0.4205(07) $188\times\ \ \!3$
E28P35 6.308 0.08973(20)(53) --0.2490 --0.2170 $28^{3}\times 64$ 351.4(1.4) 717.94(93) 4.462(17) --0.2201(3) 0.2823(25) $147\times\ \ \!4$
F32P30 6.410 0.07753(03)(45) --0.2295 --0.2050 $32^{3}\times 96$ 300.4(1.2) 675.98(97) 3.780(15) --0.2038(1) 0.1974(05) $250\times\ \ \!3$
F48P30 --0.2295 --0.2050 $48^{3}\times 96$ 302.7(0.9) 674.76(58) 5.713(16) --0.2037(1) 0.1965(04) $\ \ \!99\times\ \ \!3$
F32P21 --0.2320 --0.2050 $32^{3}\times 64$ 210.3(2.3) 658.79(94) 2.646(28) --0.2023(1) 0.1996(04) $194\times\ \ \!3$
F48P21 --0.2320 --0.2050 $48^{3}\times 96$ 207.5(1.1) 661.94(64) 3.917(22) --0.2025(1) 0.1997(04) $\ \ \!82\times 12$
G36P29 6.498 0.06887(12)(41) --0.2150 --0.1926 $36^{3}\times 108$ 297.2(0.9) 693.05(46) 3.731(11) --0.1928(1) 0.1433(12) $270\times\ \ \!4$
H48P32 6.720 0.05199(08)(31) --0.1850 --0.1700 $48^{3}\times 144$ 316.6(1.0) 691.88(65) 4.000(12) --0.1701(1) 0.0551(07) $157\times 12$
Lattice QCD discretizes the fermion and gauge actions on the lattice using dimensionless input parameters $\tilde{m}^{\rm b}_{q}$ for the masses of different quark flavors, as well as the effective gauge coupling $\hat{\beta}=10/(g^{2})$. Here, we use $\tilde{O}$ to denote the dimensionless value of any quantity $O$. After the fermion fields are integrated out, their impact can be expressed as the determinants of the D-slash operator as a function of the gauge field and $\tilde{m}^{\rm b}_{q}$. The functional integration of the gauge field on a lattice with dimensionless volume $\tilde{L}^{3}\times\tilde{T}$ is approximated by statistical samples generated through the hybrid Monte Carlo (HMC) algorithm Duane et al. (1987). The physical scale of QCD emerges when we calculate the dimensional gauge quantities, which are proportional to $\Lambda_{\text{QCD}}^{n}$, such as the scale parameter $w_{0}$=0.1736(9) fm Aoki et al. (2022); Borsanyi et al. (2012); Bazavov et al. (2014); Blum et al. (2016) for the system with $N_{f}=2+1$. The lattice spacing $a$ can then be obtained from the ratio of $w_{0}$ to its dimensionless value $\tilde{w}_{0}$.
The results in this work, are based on the 2+1 flavor (degenerate up and down quarks plus the strange quark) ensembles Hu et al. (2024) from the CLQCD collaboration using the tadpole improved tree level Symanzik (TITLS) gauge action and the tadpole improved tree level Clover (TITLC) fermion action. As discussed in the recent lattice average of flavor physics Aoki et al. (2022), the systematic uncertainty stemming from the absence of the charm quark as a dynamical degree of freedom in the action is expected to vanish within the uncertainty due to the decoupling of the heavy quark with $m_{Q}\gg\Lambda_{\rm QCD}$. More details of the gauge ensembles and verification of the independence between statistical samples of all the ensembles can be found in Sec. A of the supplemental materials.
@@
Figure 1: The $m_{\eta_{s}}^{2}$ dependence of the $D_{s}^{*}$ mass at $a=0.1053$ fm, with sea strange quark mass parameter (red dots) and valence one (blue crosses).
The parameters employed for the simulation are outlined in Table 1. The values of $m_{\pi}$ of all the ensembles except F32P21 are chosen to ensure $m_{\pi}L\geq$3.7, thereby efficiently suppressing the finite volume effect. The “fictitious” meson $\eta_{s}$ corresponds to the octet pseudoscalar meson with strange quark mass, which can be precisely determined on the lattice once the physical strange quark mass is derived from other physical quantities, such as $m_{K}$ or $m_{\Omega}$. For instance, we obtained $m_{\eta_{s}}=687.4(2.2)\,\text{MeV}$ using the physical strange quark mass derived from $m_{K}$ Hu et al. (2024) on the CLQCD ensembles.
Based on the framework of partially quenched QCD, the mass of the quark associated with the hadron interpolation field—referred to as the “valence” mass—can differ from that in the gauge configuration, which corresponds to that in the “sea” quark loop. Usually the impact of the valence quark mass is much stronger than that of the sea quark mass. Thus one can tune the valence strange quark mass by the most precise lattice determination $m_{\eta_{s}}=$ 689.89(49) MeV Borsanyi et al. (2021) on each ensemble to suppress the impact of unphysical strange quark mass on the physical observables.
@@
Fig. 1 illustrates the strange quark mass dependence (through $m_{\eta_{s}}^{2}$) of the $D_{s}^{*}$ meson mass at $a=0.1053$ fm and the tuned valence charm quark mass, using either the sea strange quark mass parameter $\tilde{m}_{s}$ (red circles) or the valence one $\tilde{m}^{\rm v}_{s}$ (blue crosses). The strange quark mass dependence of the latter case is only 10% of the former case and then allows a more reliable continuum extrapolation since the contribution of the cross term likes $m_{\eta_{s}}^{2}a^{2}$ will be highly suppressed. The residual deviation between the sea strange quark mass and the physical one will be addressed through a joint fit using all the ensembles.
@@
Determining the valence charm quark mass can present additional complexities. The physical $J/\psi$ or $\eta_{c}$ mass should account for both the disconnected charm quark contribution and QED effect, making it unsuitable for accurately determining the charm quark mass using pure QCD calculations with only the connected charm quark contribution. Conversely, the open-charm pseudoscalar meson masses are free of the disconnected valence charm quark contribution, and its QED effect has been computed in previous literature Giusti et al. (2017) and updated in Ref. Di Carlo et al. (2019). Compared to the strange quark, the valence light quark with physical mass is noisy, costly, and also susceptible to significant finite-volume effects on most of our ensembles. Thus the optimal choice to determine the interpolated valence charm quark mass parameter $\tilde{m}^{\rm v}_{c}$ on each ensemble, is requiring the $D_{s}$ meson mass to be the QED effect subtracted physical value Di Carlo et al. (2019) under the GRS renormalization scheme $m^{\overline{\textrm{MS}}}_{q,\rm QCD+QED}(2\mathrm{GeV})=m^{\overline{\textrm{MS}}}_{q,\rm QCD}(2\mathrm{GeV})$ Gasser et al. (2003),
@@
Since the mass $m_{D_{s}}$ remains unchanged when we interchange the up and down quark masses in pure QCD, all terms of the form $(m_{d}-m_{u})^{2n+1}$ will vanish. The next leading order isospin-breaking effects will thus be of the order ${\cal O}((m_{d}-m_{u})\alpha)$ and ${\cal O}(\alpha^{2})$, which are negligible in comparison to the leading order QED effects.
The values of the bare valence strange and charm quark masses $\tilde{m}^{\rm v}_{s,c}$ are also presented in Table 1, based on the polynomial interpolation of the calculations of $m_{\eta_{s},D_{s}}$ using three strange (charm) mass parameters.
We observe that $\tilde{m}^{\rm v}_{s,c}$ on ensembles with the same lattice spacing exhibits slight variations, which can be attributed to the differences in the light and strange sea quark masses.
@@
where $A_{\mu}=\bar{\psi}\gamma_{5}\gamma_{\mu}\psi$, $P=\bar{\psi}\gamma_{5}\psi$, and $m_{\rm PS}$ is the corresponding pseudoscalar meson mass. The renormalized quark mass can be subsequently defined as $m^{R}_{q}=Z_{A}/Z_{P}m^{\rm PC}_{q}$, where $Z_{A,P}$ are the renormalization constant of corresponding currents.
Following the previous light quark study on the same ensembles, we use the coulomb wall source propagator to suppress the statistical uncertainty and extract the meson mass and local operator matrix elements through the joint fit of the wall-to-point and wall-to-wall two-point functions Hu et al. (2024). The number of the configuration $n_{\rm cfg}$ and sources used per configuration $n_{\rm src}$ are also collected in Table 1. In the case of charmed hadrons, a multi-state fit is essential to accurately determine the ground state mass. As an illustration, we present our fitting results on the H48P32 ensemble, which features the finest lattice spacing, in Sec. B of the supplementary materials. Consistency check on two determinations of $m_{s}^{\rm PC}+m_{c}^{\rm PC}$ from quarkonium ($\eta_{s}$ and $\eta_{c}$) and also $D_{s}$ can be found there.
Then the meson decay constants can be extracted from the matrix elements with proper normalization,
@@
All renormalization constants can be obtained by taking their ratio over $Z_{V}$, once $Z_{V}$ is determined. But unlike its continuum counterpart, $Z_{V}$ under lattice regularization is subject to both $\alpha_{s}$ and $a^{2}$ corrections and can be determined from the vector current conservation condition Martinelli et al. (1995); Zhang et al. (2021),